OpenStudy (anonymous):

Trish has a jug of contaminated water that she knows is only 75% clean. She also knows that the water will be safe to drink if she can mix it with 100% clean water so that it is 97% clean. How much clean water (100% clean) would Trish need to mix with the contaminated water in order to get 1 gallon of drinkable water?” A student used the following system of equations to model the situation given above. Which of the statements is true about the system of equations the student used? (75x+100y=97(1) x-y=1

OpenStudy (anonymous):

system of equation is missing

OpenStudy (anonymous):

ok im bout to put it on there for you

OpenStudy (anonymous):

this is the choices A. The system of equations the student used is correct. B. The top equation in the system is incorrect. C. The bottom equation in the system is incorrect. D. Both equations in the system are incorrect.

OpenStudy (anonymous):

1) 75x+100y=97/(x+y) eq was incorrect 2) x+y=1 gallon=97/(75x+100y) eq was incorrect no guarantee

OpenStudy (anonymous):

thank you soo much